Please provide a descriptive explanation behind the work used to solve the problemProblem 11-Imaging: You are given a plano-convex lens. The glass has a refractive index of 1.5. The magnitude of radius of cu
Answer #1

soll 11 n=1.5 Ri= 100mm di = 2m - Ooloom Griven Data Part (a) Using the relation of focal length 1-6-16-1) {R-00] 1/2 = (lis-U - (2 +U) (2+UU 1 200 x103 -2 2u+U2 200*1003 2u+U? + 400*10-3 U2+2uto.t=0 Tua U= -0.2254m),–17746 m x VE 2-0.2254 m the lone